OpenStudy (anonymous):

Simplify. cos5π/12 cosπ/12-sin5π/12 sinπ/12

OpenStudy (anonymous):

this is the addition angle formula, but this time for cosine

OpenStudy (anonymous):

making it \[\cos(\frac{5\pi}{12}+\frac{\pi}{12})=\cos(\frac{\pi}{2})\]